OpenStudy (anonymous):

multiply 9.041*10^2 by .082*10^-1

OpenStudy (anonymous):

is it 74135.2?

OpenStudy (anonymous):

9.041 x 10^2 = 904.1 .082 x 10^-1 = .0082 Can you multiply it now?

OpenStudy (unklerhaukus):

\[(9.041\times10^2)\times (.082\times10^{-1})\] \[=(9.041\times.082)\times (10^2\times10^{-1})\]